# AutoScale: Automatic Prediction of Compute-optimal Data Composition for Training LLMs

### v2

@@ 2 Related Work
Domain Reweighting. An emerging line of research strives to optimize the composition of training data for LLMs pre-training with domain reweighting , i.e., adjusting the relative proportion of data from different data sources to "best" (in terms of training efficiency, final model performance, etc.) train the model. DoReMi (Xie et al., 2024) first trains a small reference model, and then trains a second proxy model with GroupDRO (Sagawa et al., 2019) to minimize the excessive domain loss relative to the reference model, where the domain weights of the proxy model will be the output. DOGE (Fan et al., 2023) trains a proxy model while tracking the first-order gradient of the model on evaluation domains (i.e., data influence) and optimizes domain weights based on the gradients, relying on infinitesimal approximations which may or may not be accurate for models trained with a practical learning rate. Data Mixing Laws (Ye et al., 2024) trains a number of proxy models to run a coarse grid search on the space of data mixtures and interpolate their performance with exponential functions to find the minimum. Similarly, RegMix(Liu et al., 2024) trains a regression model to represent the relationship between training data mixtures and resulting model performance and optimize data composition based on it.
These methods often rely on ad-hoc hyperparameter tuning via trial and error, achieving varying results. Further, the optimized weights are directly applied to training the target model on magnitudes of larger data scales. This implicitly poses a strong assumption that the "optimal data composition" is invariant of model sizes or data scales. Yet, optimal data composition is likely to shift with data size. Optimal curation at a smaller scale may not remain optimal at the target scale (Sorscher et al., 2022; Goyal et al., ). (Albalak et al., 2024) provides a recent survey for this fast-evolving field. We refer to App. A for broader discussions.
Scaling Laws. Extensive research shows that Neural Scaling Laws, predicting how the model performance changes with the scale of training data, model parameters, and computation budget (Kaplan et al., 2020), to be accurate in various tasks from vision and text processing (Alabdulmohsin et al., 2022) to LLM pre-training (Rae et al., 2021) and evaluations (Gadre et al., 2024). (Hoffmann et al., 2022) proposes compute-optimal scaling for LLM pretraining data scales together with the model’s parameter sizes. Yet, recent progress (AI@Meta, 2024; Mehta et al., 2024) shows no sign of saturation in pre-training even for models pre-trained on a considerably larger data scale than recommended by (Hoffmann et al., 2022). (Bahri et al., 2021) shows that data from different sources generally scale at different rates. Seminal work (Sorscher et al., 2022) sheds light on the possibility of attaining beyond-neural scaling law performance if one could find the best training dataset for each training data scale. This work is connected to the research on scaling laws in two ways. First, we leverage scaling laws to model the functional relationship between the quantity of data from each domain and trained model performance, allowing optimizing the training data composition in a reasonable time with high precision; further, this work contribute to a novel dimension of scaling laws–scaling optimal data compositions with the training data scale, providing original insights, clear empirical evidence, and theoretical frameworks which enable further analysis.
3 Optimal Data Composition is Scale-Dependent and Predictable
For "compute-optimal" domain weights, the goal is to find an optimal training data composition such that, for a given compute budget (i.e., training data size), the empirical validation loss, measured in perplexity (PPL), is minimized (Xie et al., 2024; Albalak et al., 2023; Fan et al., 2023). Formulating this as a bi-level optimization problem, in this section, we first introduce an original solution approach via scaling law approximations, which allows solving it efficiently and effectively. Then, with this solution approach, we solve for the optimal domain weights under different training data scales. Our results demonstrate that the optimal data composition for a fixed compute budget depends on the scale of the training data. Via the lens of scaling laws, this work pioneers in deriving an analytical framework for modeling the functional relationship between optimal data composition and training data scales.
3.1 Compute-Optimal Training Data Compositions
Consider training an LLM on a data composition $S$ from $m$ domains, $D_{1},D_{2},\cdots,D_{m}$. Let $S=\{S_{1},S_{2},\cdots,S_{m}\}$ denote the training dataset where $S_{i}$ is the subset of training data from each domain. The domain weights $\mathbf{w}=[w_{1},w_{2},\cdots,w_{m}]^{T}$ are defined as the proportions of data for each domain. Namely, letting $N=|S|$ denote the amount of total tokens of training data, domain weights are given as $w_{i}=N_{i}/N$, where $|S_{i}|$ denotes the amount of tokens for training subset $S_{i}$.
Let $\bm{\theta}^{*}(S)$ denote the parameters of a learning algorithm (i.e., the model) trained on data $S$ with empirical risk minimization (ERM), given as $\bm{\theta}^{*}(S):={\arg\min}_{\bm{\theta}}\mathcal{L}(\bm{\theta},S)$ where $\mathcal{L}(\bm{\theta},S)$ denotes the loss of model parameterized by $\bm{\theta}$ evaluated on data $S$, which is the training loss. Since training data $S$ can be equivalently defined by its data quantity and domain weights $(N,\mathbf{w})$, we define a slight change of notation $\bm{\theta}^{*}(N,\mathbf{w}):=\bm{\theta}^{*}(S)$ and will use $S$ and $(N,\mathbf{w})$ interchangeably. We would like to maximize the amount of information gain and achieve maximum loss reduction during training, given as $\min_{\mathbf{w}\in\mathbb{W}^{m}}\mathcal{L}(\bm{\theta}^{*}(N,\mathbf{w}),D^{v})=\min_{\mathbf{w}\in\mathbb{W}^{m}}\sum_{i=1}^{m}\mathcal{L}(\bm{\theta}^{*}(N,\mathbf{w}),D_{i}^{v})$, where $D^{v}$ and $D^{v}_{i}$ denote total validation data and validation data of individual domain $i$, respectively; the space of weights $\mathbb{W}^{m}$ is the hyperplane of the probability simplex $\mathbb{W}^{m}=\{\mathbf{w}|w_{1}+w_{2}+\cdots+w_{m}=1\}\cap\{\mathbf{w}|0\leq w_{i}\leq 1,\forall i\in\{1,2,\cdots,m\}\}$. Define minor simplifications of notations for the validation losses $\mathcal{L}^{v}(\theta,D^{v}):=\mathcal{L}(\theta,D^{v})$ and $\mathcal{L}^{v}_{i}(\theta,D^{v}):=\mathcal{L}(\theta,D^{v}_{i})$. Then, the optimal domain weights, $\mathbf{w^{*}}$, are given as the minimizer of the objective,
$\mathbf{w^{*}}=\arg\min_{\mathbf{w}\in\mathbb{W}^{m}}\sum_{i=1}^{m}\mathcal{L}_{i}^{v}(\bm{\theta}^{*}(N,\mathbf{w}))\quad\text{s.t.}\,\,\,\,\bm{\theta}^{*}(N,\mathbf{w})=\arg\min_{\bm{\theta}}\mathcal{L}({\bm{\theta}},(N,\mathbf{w}))\vspace{-1.3em}$ (1)
where perplexity is adopted as the loss metric. This formulation is a bi-level optimization problem, where the outer problem seeks the optimal domain weights, while the inner problem is training the model with ERM on the data defined by certain weights. A general approach is to solve it with gradient descent, $\mathbf{w}^{t+1}=\mathbf{w}^{t}-\eta\cdot\frac{\partial\mathcal{L}_{v}(\bm{\theta}^{*}(N,\mathbf{w}^{t}))}{\partial\mathbf{w}}$. Since there is no tractable form of analytical expression for $\bm{\theta}^{*}$, this gradient needs to be estimated with empirical methods (e.g., approximated by finite difference), requiring repetitive re-training of the model at each update Liu et al. (2021).
3.1.1 Solution via Scaling-law-inspired Approximations
Directly optimizing training data composition by solving bi-level optimization problems involves repetitive model retraining, which can be prohibitively expensive even at small scales. Current work (Xie et al., 2024; Fan et al., 2023; Ye et al., 2024; Liu et al., 2024) mostly employs heuristic methods to conduct this optimization on smaller models trained with fewer data, achieving varying results in different cases. To crystalize the relationship between optimal data compositions and training data scales and obtain a clear image of the complete landscape, we propose an original approach to this problem. We propose to first fit a scaling function to the outer loss (validation loss) $\mathcal{L}^{v}$ as a function of domain weights $\mathbf{w}$, effectively reducing the bi-level problem to a single level, which can be solved efficiently via regular gradient descent, allowing finding the global optimum efficiently and accurately.
Figure 2: Fitting validation loss with power-law functions for 774M Decoder-only LMs (GPT-2 Large), directly approximating how loss changes with each domain’s data quantity.
To begin with, neural scaling laws suggest the relationship between a model’s evaluation loss and the size of its training data can be well-represented by power law functions (Kaplan et al., 2020) $\mathcal{L}_{v}(\bm{\theta}^{*}(N,\mathbf{w}))=N^{-\gamma}+\ell_{0}$ where constants $\ell_{0}$ denotes some irreducible loss and $\gamma\geq 0$ is some scaling coefficient. Drawing inspirations from (Hernandez et al., 2021), which formulates the scaling laws for transfer learning, we propose the following approximation to model the scaling relationship between model loss and training data quantity from different sources/domains.
Consider a model trained on data with size $N$ and domain weights $\mathbf{w}$. Define constant $N_{0}^{i}$ which estimates the evaluation loss when the amount of training data from domain $i$ is zero (i.e., $N_{i}^{\prime}=0$), which effectively measures the effect of data from all other domains. From this regard, $N_{0}^{i}$ can be interpreted as the equivalent data size for training data from domains other than $i$. Notably, this formulation aligns with empirical findings in the prior literature (Hernandez et al., 2021; Ye et al., 2024). Then, for training data defined by $(N,\mathbf{w})$ where the amount of training data from domain $D_{i}$ is $N_{i}=N\cdot w_{i}$, evaluation loss can be expressed as a function of $N_{i}$: $\mathcal{L}_{v}(\bm{\theta}^{*}(N,\mathbf{w}))=(N_{0}^{i}+N_{i})^{-\gamma_{i}}+\ell_{i}$ where $\gamma_{i},\ell_{i}$ are constants associated with domain $i$. If the amount of training data from one domain $D_{i}$ is changed from $N_{i}$ to $N_{i}^{\prime}$ with the amount of training from other domains unchanged, we approximate the new model’s evaluation loss, $\mathcal{L}_{i}^{\prime}$, after re-training with a power law function of $N_{i}^{\prime}$:
$\vspace{-1.5em}\mathcal{L}_{v}(\bm{\theta}^{*}(N^{\prime},\mathbf{w^{\prime}}))=(N_{0}^{i}+N_{i}^{\prime})^{-\gamma_{i}}+\ell_{i}:=\mathcal{L}_{i}^{\prime}\vspace{-1em}$ (2)
where $N^{\prime}=N+(N_{i}^{\prime}-N_{i})$ denotes the updated amount of training data, and $w_{i}^{\prime}=N_{i}^{\prime}/N^{\prime}$ denotes the updated domain weights.
We propose the following procedure to fit the parameters in Eq. (2). We re-train two models with different data quantities for domain $i$, $N_{i}^{+}$ and $N_{i}^{-}$ where $N_{i}^{-}<N_{i}<N_{i}^{+}$ [^2] Empirically, we found setting the perturbation ratio, $r=N_{i}/N_{i}^{-}=N_{i}^{+}/N_{i}=3$, produces reliable results., and compute their evaluation loss, $\mathcal{L}_{i}^{+}$ and $\mathcal{L}_{i}^{-}$, respectively. Then, together with evaluation loss $\mathcal{L}_{v}^{0}=\mathcal{L}_{v}(\bm{\theta}^{*}(N,\mathbf{w}))$ for the original model trained with $N_{i}$, the parameters $\gamma_{i}$, $\ell_{i}$ and $N_{0}^{i}$ can be estimated via ordinary least square (OLS) fitting,
[…]

### v5

@@ 2 Related Work
Principled training data curation for LLMs is an emerging research area, aiming to strategically select data that improves model performance. It can be performed at multiple levels—from token-level (Lin et al., 2024) or point-level (Wang et al., 2024) up to domain-level selection. Domain-level approaches are often more efficient because they operate at a coarser granularity, typically applying soft selection—i.e., upweighting or downweighting entire data domains.
Domain reweighting can generally be viewed as a two-step process: (i) define an objective that captures the goal of improving test loss or other model performance measures, and (ii) optimize domain weights according to that objective. DoReMi (Xie et al., 2024), a seminal paper in the space, adopted GroupDRO (Sagawa et al., 2019) as the objective, which implicitly upweights “difficult” domains. However, subsequent studies (Fan et al., 2023) found that the performance gains are unstable and limited, partly because the chosen objective does not align with the metrics ultimately used to evaluate the model at test time. Recent methods (Liu et al., 2024; Ye et al., 2024; Fan et al., 2023) attempt to directly optimize validation loss, which serves as a closer proxy for the test-time model performance metrics we actually care about. We note that there can still be misalignments between validation loss and test-time performance metrics (Barton, 2024), which is an active research area in itself, but validation loss remains a widely accepted objective for model selection in pre-training.
With validation loss as the objective, the next challenge is how to optimize it. Evaluating the objective for a given set of weights is computationally expensive, as it requires training a model from scratch on the weighted data set and evaluating the corresponding validation loss. Existing approaches tackle this in two broad ways. One line uses surrogate models to approximate the mapping from domain weights to model performance (Ye et al., 2024; Liu et al., 2024); fitting such models can still require large amounts of retrainings. For instance, in Liu et al. (2024), fitting the surrogate requires more than ten times as many training runs as there are domains. Another line performs local approximations, assuming only a single gradient step for the underlying model (Fan et al., 2023), which may not hold for practical learning rates.
Our work follows the surrogate-modeling line but differs by proposing a new parametric function to model performance versus domain weights, which can be reliably fit with only about twice as many retraining runs as the number of domains. Crucially, existing methods often conduct this optimization at a smaller data scale to keep cost manageable, then directly apply the small-scale “best” mixture at large scale. We show that this scale-invariance assumption can break down in practice. To address this, we contribute a novel theoretical analysis that characterizes how the optimal domain ratio shifts across scales, enabling us to extend small-scale insights effectively to larger budgets.
Finally, the general notion that data curation should be scale-dependent has appeared in prior works (Sorscher et al., 2022; Goyal et al., ). However, Sorscher et al. (2022) argues this point using a simplified analysis with a perception model—showing that larger scales favor “harder” samples while smaller scales favor “easier” samples—but does not propose a practical pipeline for scale-aware data selection. Our work addresses this gap by introducing a concrete method for scale-dependent curation at foundation-model scale. Meanwhile, Goyal et al. focuses on the CLIP model (Radford et al., 2021), finding that different training epochs call for different data-selection thresholds. By contrast, our setting involves (i) LLM pre-training (with typically one epoch), and (ii) different data modality and selection granularity.
3 Methodology
Figure 1: Domain weights that excel at one scale may underperform at another. Weights $w_{1}$ and $w_{2}$ are obtained by running DDO (as introduced in Section 3.2.1) at 0.3B and 1.2B, respectively.
Evidence of scale dependence.
A simple experiment illustrates that domain weighting is not one‐size‐fits‐all: we derive two data mixes (with a procedure introduced later) and compare them at different training budgets. As shown in Figure 1, when tested at 0.3B tokens, Mix A beats Mix B as measured by validation perplexity reduction compared against uniform weights, but at 1.2 B tokens Mix B outperform Mix A. This flip indicates that a domain which helps more at a smaller scale may not remain a better choice at a larger scale, while another domain initially less impactful can become more valuable as training grows. Consequently, a scale‐aware approach is needed so that domain weights adapt as more tokens are introduced.
3.1 Problem Formulation
To capture how domain importance shifts with the total training budget, we now formalize the scale‐dependent domain reweighting problem. This framework lets us solve for a better domain mixture at any given data scale.
Notations and setup.
We consider $m$ domains $\{D_{1},\dots,D_{m}\}$, each with a large pool of training examples. A domain mix is specified by a weight vector $\mathbf{w}=[w_{1},\dots,w_{m}]^{\top}$ on the probability simplex $\mathbb{W}^{m}\;:=\;\bigl\{\mathbf{w}\in\mathbb{R}^{m}\,\big|\,\sum_{i=1}^{m}w_{i}=1,\,w_{i}\geq 0\text{ for all }i\bigr\}$. Given a total budget of $N$ tokens, let $N_{i}=\lfloor w_{i}\cdot N\rfloor$ be the number of tokens chosen from domain $D_{i}$. We denote this resulting dataset by $S(N,\mathbf{w})\;=\;\{S_{1},\dots,S_{m}\},$ where $S_{i}\subseteq D_{i},\lvert S_{i}\rvert=N_{i}$. Training a model $\boldsymbol{\theta}$ on $S(N,\mathbf{w})$ means solving an empirical risk minimization (ERM) objective: $\boldsymbol{\theta}^{*}(N,\mathbf{w})\;=\;\arg\min_{\boldsymbol{\theta}}\;\mathcal{L}\bigl(\boldsymbol{\theta},\,S(N,\mathbf{w})\bigr),$ where $\mathcal{L}$ is a next-token prediction loss.
Objective function.
We assess a domain mix $\mathbf{w}$ by measuring the validation loss $\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr)\;=\;\mathcal{L}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w}),\,D^{v}\bigr)$ on a held-out dataset $D^{v}$. We then seek the mix $\mathbf{w}$ that minimizes this validation metric at scale $N$:
$\mathbf{w}^{*}\;=\;\arg\min_{\mathbf{w}\in\mathbb{W}^{m}}\;\;\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr).$ (1)
Because $\boldsymbol{\theta}^{*}$ depends on $\mathbf{w}$ via ERM, this becomes a bi-level optimization problem. Since no closed-form expression exists for $\boldsymbol{\theta}^{*}$, any gradient-based approach must approximate $\tfrac{\partial}{\partial\mathbf{w}}\,\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr)$. Traditional bi-level methods rely on higher-order derivatives with respect to model parameters for this approximation, but such techniques become infeasible at the scale of modern foundation models (Liu et al., 2021).
3.2 Our Solution
We propose a two-stage framework, AutoScale, for finding scale‐aware data compositions, which first approximates the optimal data mix at small scales and extrapolates to a larger target scale:
  1. 1.
    Direct Data Optimization (DDO): At smaller, computationally feasible scales, we learn a mapping from domain weights to validation loss. This reduces the original bi-level problem to a single-level convex optimization—letting us approximate the “best” domain mix for that smaller budget.
  2. 2.
    Optimal Mix Projection: Building on a theoretical analysis of how domain importance changes with total tokens, we then extrapolate those small‐scale DDO solutions to a larger data budget.
3.2.1 Direct data optimization
Direct Data Optimization (DDO) is a practical method for approximating the solution to the bi-level domain‐weighting problem at relatively small data scales. The key idea of DDO is to approximate the validation loss $\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr)$ as a parametric function of the domain‐weight vector $\mathbf{w}$. This effectively reduces our bi‐level objective (choose $\mathbf{w}$ while also training $\theta$) to a single‐level optimization, which can be solved efficiently via standard gradient‐based methods.
We begin by noting that the validation loss can be represented by scaling laws as a function of data size for each individual domain. We model the dependence of validation loss on the size of data from each domain, then aggregate these functions to derive the final approximation for the validation loss on $\mathbf{w}$.
Drawing inspiration from neural scaling laws—which indicate a power-law relationship between training data scale and validation loss (Kaplan et al., 2020)—we assume that validation loss as a function of domain $i$’s data size follows
$\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr)\;\approx\;\bigl(N_{0}^{i}+w_{i}\cdot N\bigr)^{-\gamma_{i}}\;+\;\ell_{i}.$
Here, $w_{i}$ denotes the fraction of the total token budget $N$ allocated to domain $i$. The term $N_{0}^{i}$ represents an the “equivalent data size” contributed by domains other than $i$, while $\gamma_{i}$ governs how quickly domain $i$ reaches a point of diminishing returns. Lastly, $\ell_{i}$ represents the irreducible term in the loss function.
To learn these parameters $\{N_{0}^{i},\gamma_{i},\ell_{i}\}$ for each domain $i$, we retrain the model after perturbing $w_{i}$ upward and downward, measure the change in total validation loss, and then fit $(N_{0}^{i}+w_{i}\cdot N)^{-\gamma_{i}}+\ell_{i}$ via least squares. Because $\mathcal{L}^{v}$ aggregates the effects of data size for each domain, our final approximation for the validation loss is:
$\mathcal{L}^{v}\bigl(\boldsymbol{\theta}^{*}(N,\mathbf{w})\bigr)\;\approx\;\sum_{i=1}^{m}\Bigl(N_{0}^{i}+w_{i}\cdot N\Bigr)^{-\gamma_{i}}\;+\;\ell_{i}.$
Once we have fitted the parameters, we can directly optimize over $\mathbf{w}$ subject to $\sum_{i=1}^{m}w_{i}=1$ to approximate the optimal domain mix under the total token budget $N$.
Because DDO only requires retraining at $(2m+1)$ mixes (one baseline plus up/down perturbations for each of the $m$ domains), it is far cheaper than a naive zero-order method that retrains the model at every weight update. Nevertheless, DDO is best suited for moderate domain counts ($m$) and data scales ($N$). For much larger target scales, we introduce a second stage that extrapolates the “best” DDO mix from smaller scales to significantly bigger budgets, all without additional retraining.
3.2.2 Optimal mix projection
Our method for extrapolating domain mixes to much larger training budgets hinges on a novel theoretical result that we developed to characterize how the optimal mix ratio depends on the total data scale. Note that all inverses $(\cdot)^{-1}$, products, and exponentiations on vectors below are understood elementwise.
Theorem 1: Scale-Dependent Optimal Composition Consider the optimization problem $\min_{\mathbf{N}}\biggl\{\sum_{i=1}^{m}\beta_{i}\,N_{i}^{-\gamma_{i}}\;\Big|\;\sum_{i=1}^{m}N_{i}=N\biggr\},$ where $\beta_{i}\geq 0$ and $\gamma_{i}\geq 0$ for all $i$, and $\mathbf{N}=(N_{1},\dots,N_{m})$ denotes the domain allocations. Let $\mathbf{N}^{*}(N)$ be the optimal allocation that minimizes the sum above for a total budget $N$. For two distinct budgets $N^{(1)}\neq N^{(2)}$, and any larger budget $N^{(3)}$, suppose there is a constant $k>0$ such that $\mathbf{N}\bigl(N^{(3)}\bigr)\;=\;\mathbf{N}^{*}\bigl(N^{(2)}\bigr)\;\Bigl[\bigl(\mathbf{N}^{*}\bigl(N^{(1)}\bigr)\bigr)^{-1}\;\mathbf{N}^{*}\bigl(N^{(2)}\bigr)\Bigr]^{k},\quad\text{with }\sum_{i=1}^{m}N_{i}\bigl(N^{(3)}\bigr)=N^{(3)}.$ Then $\mathbf{N}\bigl(N^{(3)}\bigr)$ is also the optimal allocation for the budget $N^{(3)}$, i.e., $\mathbf{N}\bigl(N^{(3)}\bigr)\;=\;\arg\min_{\mathbf{N}}\Bigl\{\sum_{i=1}^{m}\beta_{i}\,N_{i}^{-\gamma_{i}}\;\Big|\;\sum_{i=1}^{m}N_{i}=N^{(3)}\Bigr\}\;=\;\mathbf{N}^{*}\bigl(N^{(3)}\bigr).$
Proof overview (high-level).
At optimality, the first-order (KKT) conditions impose that each domain’s partial derivative of the loss matches up to a single Lagrange multiplier. From this, we can derive how each domain’s optimal allocation $N_{i}^{*}(N)$ scales when transitioning from budget $N^{(1)}$ to another budget $N^{(2)}$. These domain-by-domain scaling factors do not depend on the absolute size of $N$, only on the relative shifts between domains, which in turn yields an exponential-style expression for the optimal allocation at a third budget $N^{(3)}$. Thus, once we know the optimal allocations at two budgets, we can directly construct the optimal allocation for any larger budget without re-solving the entire optimization.
Interpretation of the theory.
The statement above assumes each domain $D_{i}$ contributes $\beta_{i}\,N_{i}^{-\gamma_{i}}$ independently to the total validation loss. In Appendix B.2, we generalize this to cases where domains may overlap by treating the evaluation as composed of multiple “latent skills” (Tiong et al., ); the same exponential-style scaling behavior still emerges.
We defer the full proof to Appendix B.1 (where we employ first-order optimality/KKT conditions), but the key insight is that domains saturate at different rates depending on their exponents $\gamma_{i}$. Specifically, a domain $D_{i}$ with a small $\gamma_{i}$ saturates more slowly and thus continues to yield benefits at larger budgets, receiving an increasingly bigger fraction of tokens as $N$ grows. In contrast, a large $\gamma_{i}$ indicates that $D_{i}$ quickly saturates, so it is favored primarily at smaller scales.
Concretely, this means the “optimal mix ratio” is not constant across all scales. As the total budget $N$ increases, domains with smaller $\gamma_{i}$ are allocated a larger share. The theorem’s exponential-style update precisely captures these changing allocations, enabling us to predict the best mix at a higher budget given the solutions at two smaller budgets—without re-solving the entire optimization problem.
3.2.3 Overall algorithm
Having established that the optimal domain mix varies predictably with training budget, we now summarize AutoScale , our proposed two-stage approach to optimize data mix.
Stage 1 (pseudocode provided in Algorithm 1): Pick two feasible scales $N^{(1)}$ and $N^{(2)}$ (with $N^{(1)}<N^{(2)}$), where retraining the model is still affordable. Run DDO to obtain optimal allocations $\mathbf{N}^{*}(N^{(1)})$ and $\mathbf{N}^{*}(N^{(2)})$.
Stage 2 (pseudocode provided in Algorithm 2): Leveraging our theoretical result, we automatically predict the optimal domain mix for any larger scale. Specifically, starting from the optimal domain allocation $\mathbf{N}^{*}\bigl(N^{(2)}\bigr)$, we repeatedly “scale up” by multiplying by $\Bigl[(\mathbf{N}^{*}(N^{(1)}))^{-1}\,\mathbf{N}^{*}\bigl(N^{(2)}\bigr)\Bigr]^{\delta}$. Each such update yields a new allocation at a larger budget than before. We continue until reaching or exceeding the target budget $N^{\text{tgt}}$. By adjusting the resolution $\delta$, we control the granularity of these updates, thus reaching the target scale with any desired accuracy.
4 Evaluation
Our evaluation aims to address the following questions:
  • (RQ1) Does DDO yield better domain weighting at smaller scales? In our approach, DDO is designed to approximate the best data mix at a given scale. While we cannot verify its optimality, we want to see if DDO meaningfully improves domain weighting compared to baseline methods (Section 4.1).
  • (RQ2) Can AutoScale—DDO at smaller scales plus our theoretical projection to larger scales—achieve training efficiency and performance benefits when direct DDO at large scale is prohibitively expensive?(Section 4.2)
Overview of experimental settings.
We provide an overview here and defer the full details to Appendix C. (I) Models and datasets. We experiment with two architectures—GPT-2 Large (774M parameters) and BERT (110M), training on up to 10B tokens. While this budget is comparatively small for the latest LLM regimes, it already exceeds the data scales used in many existing domain-reweighting studies (Fan et al., 2023; Chen et al., 2024a), serving as a feasible testbed for prototype ideas in non-commercial settings. Specifically, we pre-train GPT-2 Large on the RedPajama dataset (Computer, 2023),which spans seven domains (e.g., Common Crawl, C4 (Raffel et al., 2020), GitHub, Wikipedia, ArXiv, StackExchange). For BERT, we use data from five sources—Amazon Reviews, Arxiv, Books, Wikipedia, and Open WebText (Gokaslan & Cohen, 2019). (II) Baselines. We compare our methods against several baseline strategies. Uniform samples data from each domain, leading to the same token count per domain. Llama weights are a curated set of heuristically tuned domain weights from the LLaMA-1/2 models (Touvron et al., 2023). DoReMi (Xie et al., 2024) is a seminal paper in this domain-reweighting space, offering an early, principled approach to finding domain weights. Data mixing law (Ye et al., 2024) and RegMix (Liu et al., 2024) represent the latest state-of-the-art. (III) Metrics. We measure test perplexity and also evaluate downstream performance to confirm that improvements extend to practical tasks.
4.1 Evaluating DDO
Effectiveness of DDO-optimized weights.
We perform DDO on GPT-2 Large at two different data scales (0.3B and 1.2B tokens) to obtain DDO weights specifically optimized for each scale. We then retrain the model under these DDO-derived weights and compare the evaluation loss against two baselines: (1) Uniform (no reweighting), and (2) RegMix, the latest state-of-the-art approach. (We omit DoReMi here, as it has been surpassed by RegMix (Liu et al., 2024).) For each set of weights, the model is trained at both 0.3B and 1.2B tokens, with results in Table 1. At both scales, DDO-optimized domain weights significantly outperform the Uniform baseline, achieving a notably lower evaluation loss. DDO-optimized weights also surpass RegMix when models are trained at the same scale as the domain-weight optimization, indicating that DDO finds more effective domain weights than RegMix. Further, RegMix does not consider adaptation for training models at different scales. Applying RegMix optimized weights on larger data scales appears less effective, evident by the widening gap between its performance from DDO’s. Notably, the DDO-derived weights yield the strongest gains at the scale for which they were optimized, while showing less advantage when used at a different scale, highlighting the scale-dependent nature of domain weighting.
Weights/Actual training scale 0.3B training tokens 1.2B training tokens
Uniform Weights 48.04 28.11
RegMix Weights (optimized at 0.3B) 46.56 (-1.48) 27.86 (-0.25)
DDO Weights (optimized at 0.3B) 46.13 (-1.91) 27.09 (-1.02)
DDO Weights (optimized at 1.2B) 46.31(-1.73) 26.97 (-1.14)
Table 1: GPT-2 Large trained with DDO optimized domain weights achieve significantly reduced test perplexity compared to with non-optimized, uniform weights, also outperforming RegMix. DDO optimized weights appear most performant at the data scale they were optimized.
In addition, we apply DDO to BERT at 0.3B tokens; the resulting model performance from the DDO-optimized weights is shown in Fig. 2. These weights reduce the model’s validation loss on all training domains and on held-out non-training domains, demonstrating DDO’s effectiveness in improving training efficiency. Furthermore, when evaluated on the GLUE benchmark and the SQuAD dataset, the DDO-optimized weights also yield a notable improvement in downstream task performance.
(a) Validation Loss ($\downarrow$ lower is better)
(b) Task Performance ($\uparrow$ higher is better)
Figure 2: Optimizing domain weights with DDO algorithm for pre-training Encoder-only LMs (BERT). DDO substantially reduces validation loss. After reweighting, all training domains’ loss has decreased or remained unchanged. Out-of-domain loss on non-training domains also decreased considerably. Enhanced performance is observed on all GLUE tasks (eval metric: cola: Matt. corr., stsb: Pearson corr., rest: acc.) and SQuAD (acc.).
Analyzing DDO’s effectiveness.
Recall that the key idea of DDO is to use a power-law–based parametric function to predict validation loss from domain weights. A major factor in DDO’s effectiveness lies in the accuracy of this function. We quantify its predictive power via the average absolute relative error (AAR) between the predicted and actual losses. In our experiments, the AAR is $1.00\%$, indicating that DDO’s modeling closely reflects actual loss.
4.2 Evaluating AutoScale
Effectiveness of our extrapolated weights.
Recall that AutoScale is a two-stage pipeline: first, run DDO at smaller scales to identify domain weights, then extrapolate those weights to a larger scale. We call the resulting allocation the AutoScale weights. For GPT-2 Large, we run DDO on up to 0.6B tokens, then extrapolate to 3B and 10B tokens. Figure 3 shows the change of test perplexity during training for models trained with 10B tokens using AutoScale weights versus baseline allocations. AutoScale consistently outperforms every baseline by a 28–38% margin and also demonstrates advantageous downstream performance. Table 2 demonstrates the results on 3B tokens, revealing that AutoScale maintains its superiority in both final loss achieved and faster convergence. Table 3 examines domain-wise test perplexities; AutoScale weights significantly reduce the loss on the Books domain and improve worst-domain perplexity, also yielding a better average across domains. Finally, Table 4 evaluates eight downstream tasks. The model trained with AutoScale weights achieves the best overall performance, further underscoring the effectiveness of our extrapolated domain weights.
For BERT, we train up to 288k steps (approximately 120% of the original BERT-base budget (Devlin et al., 2018)). Table 10 shows that, compared to uniform (no reweighting), AutoScale yields a 16.7% speed-up at most data scales and a 10% speed-up at the largest scale, demonstrating consistent effectiveness. However, these gains are smaller than those observed for GPT-2 Large, indicating that different architectures and training objectives may respond differently to domain reweighting. This is also hinted at in Figure 10, where the evaluation loss shows a more uniform response to each domain, suggesting fewer benefits from reweighting in BERT’s setup.
Figure 3: Training 774M Decoder-only LMs (GPT-2 Large) for 10B tokens (96k steps). AutoScale-predicted domain weights decrease test perplexity at least $28\%$ faster than any baseline with up to $38\%$ speed up, achieving best overall task performance.
Weights Final Perplexity AutoScale Speed Improvement
(3B training tokens) (PPL) (% steps saved to final PPL)
AutoScale (ours) 21.123 -
DoReMi 21.676 25%
Data Mixing Laws 23.333 37%
LLaMA 22.944 31%
RegMix 21.740 28%
Uniform (30% more tokens) 21.736 37%
Table 2: Domain perplexity for 774M Decoder-only LMs (GPT-2 Large) trained for 3B tokens. AutoScale -predicted weights decreases val loss at least $25\%$ faster than any baseline with up to $37\%$ speed up. Despite LLaMa weights being very different from uniform weights, they yield highly similar training efficiency at these data scales.
Domain/Method AutoScale DoReMi Data Mixing LLaMA RegMix Uniform
(ours) Laws (30% more tokens)
Common Crawl 25.598 24.116 30.824 21.464 24.430 28.351
Github 7.482 6.678 5.845 7.376 6.145 5.784
Books 29.162 33.324 34.450 35.533 32.985 31.140
Wikipedia 18.828 17.154 26.795 21.110 20.177 19.570
C4 34.242 39.429 38.521 37.393 39.654 40.323
Stack Exchange 15.991 15.393 14.519 20.133 15.225 13.890
Arxiv 16.558 15.638 12.372 17.598 13.563 13.082
Average 21.123 21.676 23.333 22.944 21.740 21.736
Worst-domain 34.242 39.429 38.521 37.393 39.654 40.323
Table 3: Domain perplexity for 774M GPT-2 Large trained for 3B tokens. AutoScale  notably achieves the lowest average test perplexity while also significantly decreasing worse-domain perplexity.
Method/Task Avg pubmedqa piqa hellaswag crows_pairs boolq arc_easy truthfulqa hellaswag
(10-shot) _english _mc2 (zero-shot)
AutoScale (ours) 0.4746 0.536 0.6202 0.3021 0.5850 0.6141 0.3977 0.4385 0.3030
Uniform Weights 0.4514 0.438 0.6115 0.2923 0.5886 0.5636 0.3742 0.4526 0.2907
[…]